Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 01, 1996, pp. 1–15 Submitted: November 11, 1995. Published: April 5, 1996.

ALGORITHMS FOR DYNAMIC NEGOTIATION

Aimee Bella, Ferenc Szidarovszky, and Lucien Duckstein

Abstract.

Most conflict situations are dynamic processes where negotiation is carried out leading to that final state. In this paper, we present three techniques by generalizing the concepts of conflict and solution method so as to include dynamic features into the negotiation methodology. The three techniques are: a graph concept in a dynamic framework, a utility gain model and a negotiation algorithm. In an upper Rio Grande basin case study, we use the negotiation algorithms to resolve conflicts in water resources management.

(1991) A.M.S. Subject Classification Codes. 90,93.

Key Words and Phrases. Conflict, solution concepts, dynamic process, negotiation, graph concept, utility gained model, water allocation, Upper Rio Grande basin.

Systems and Industrial Engineering Department, University of Arizona, Tucson, AZ 85721, USA E-mail Address: szidar@sie.arizona.edu \begin{matrix}c \\ ◯ \end{matrix}1996 Cameron University

Typeset by AMS-TEX